# Invariant Consistency for Knowledge Distillation

### v1

@@ 4 Experiments
We evaluate our ICD framework in the KD task for model compression of a large network to a smaller one.
Datasets. (1) CIFAR-100 [17] contains 50,000 training images with 500 images per class and 10,000 test images. (2) STL-10 [7] consists of a training set of 5,000 labeled images from 10 classes and 100,000 unlabeled images, and a test set of 8,000 images. (3) Tiny ImageNet (TIN-200) [8] comprises 200 classes, each with 500 training images and 50 validaton images.
Setup. We experiment on CIFAR-100 with student-teacher combinations of different capacities, such as ResNet [11] or Wide ResNet (WRN) [36], VGG [28], MobileNet [26], and ShuffleNet [37, 19] (more details are described in the supplementary material). We set the temperature parameter to $\tau=0.07$. The hyperparameters are empirically set to $\alpha=0.5$ and $\beta=1$. Both the student and teacher outputs are projected to a 128-dimensional space using a projection head consisting of a single linear layer, followed by $\ell_{2}$ normalization. We train for a total of 240 epochs. More details on the training procedures can be found in the supplementary material.
@@ 4.1 Results on CIFAR-100
Table 1 and Table 2 present a comparison of top-1 accuracies for various distillation objectives. Table 1 focuses on student and teacher models with identical architectures, while Table 2 examines models with differing architectures. Our method consistently outperforms other distillation objectives, including the original KD, demonstrating an average relative improvement of 24.66% for ICD and 76.42% for ICD combined with KD. Significantly, when applying our method to distill from WRN-40-2 to WRN-16-2 (same architecture) or from WRN-40-2 to ShuffleNet-v1 (different architecture), it not only matches but surpasses the performance of the teacher network. Specifically, it achieves accuracy improvements of 0.45% and 0.78% over the teacher, respectively. ICD combined with KD performs exceptionally well even when student and teacher models have different architectures, achieving top accuracies in several configurations, such as MobileNet-v2 distilled from VGG-13 and ShuffleNet-v1 distilled from WRN-40-2. While CRD and CRD combined with KD also perform well, ICD and ICD combined with KD frequently achieve higher accuracies. We also observe that KD alone performs robustly, as no other methods except CRD and ICD consistently surpass traditional KD (as indicated by green arrows $\uparrow$ in Table 1 and Table 2). Both CRD and ICD utilize contrastive objectives to ensure consistency between the student and teacher models. Additionally, methods distilling from intermediate representations tend to underperform compared to those focusing on the last several layers when switching from same to different architectures. For instance, the Attention Transfer (AT) and FitNet methods underperform the vanilla student model, whereas PKT, SP, CRD, and ICD, which focus on the last layers, perform well. This discrepancy can be attributed to the different solution paths mapping input to output across architectures, leading to conflicts when mimicking intermediate representations.
@@
ICD 74.92 ($\uparrow$) 73.69 ($\uparrow$) 71.18 ($\uparrow$) 71.00 ($\uparrow$) 73.11 ($\uparrow$) 74.23 ($\uparrow$) 72.98 ($\uparrow$)
ICD+KD 76.06 ($\uparrow$) 74.76 ($\uparrow$) 71.81 ($\uparrow$) 71.57 ($\uparrow$) 73.62 ($\uparrow$) 74.99 ($\uparrow$) 73.83 ($\uparrow$)
Table 1: Test top-1 accuracy (%) of student networks on CIFAR-100, comparing students and teachers of the same architecture using various distillation methods. $\uparrow$ denotes outperformance over KD and $\downarrow$ denotes underperformance. Results adapted from [29]. The values in bold indicate the maximum of each column while underlined values mark the second best. The citations and abbreviations of other methods are described in the supplementary material.
@@
ICD 68.22 ($\uparrow$) 67.39 ($\uparrow$) 73.85 ($\uparrow$) 74.07 ($\uparrow$) 75.23 ($\uparrow$) 74.98 ($\uparrow$)
ICD+KD 69.37 ($\uparrow$) 69.28 ($\uparrow$) 73.88 ($\uparrow$) 75.27 ($\uparrow$) 76.53 ($\uparrow$) 76.39 ($\uparrow$)
Table 2: Test top-1 accuracy (%) of student networks on CIFAR-100 involving students and teachers from different architectures, using various distillation methods. $\uparrow$ denotes outperformance over KD and $\downarrow$ denotes underperformance. Results adapted from [29]. The values in bold indicate the maximum of each column while underlined values mark the second best. The citations and abbreviations of other methods are described in the supplementary material.
@@ 4.3 Transferability of representations
Our research focuses on transferring knowledge from a teacher network to a student network while learning representations that embody general knowledge applicable across various tasks and datasets. To investigate this, we employ a distillation process where a WRN-40-2 teacher network transfers its learned representations to a WRN-16-2 student network, which can either be trained directly from the CIFAR-100 dataset or through distillation. In our experiments, the student network functions as a fixed feature extractor, processing images from STL-10 and TIN-200, both resized to $32\times 32$. To assess the generalizability of these representations, we train a linear classifier on top of the last feature layer to perform 10-way classification for STL-10 and 200-way classification for TIN-200. The effectiveness of different distillation methods in enhancing the transferability of these representations is detailed in Table 3. Our results indicate that most distillation methods, except for FitNet, significantly enhance the transferability of the learned representations on both target datasets. Interestingly, the teacher network, while performing best on CIFAR-100, offers representations that transfer the least effectively, possibly due to a bias towards the training dataset. In contrast, the student network employing a combination of ICD and KD distillation not only matches the teacher’s performance on CIFAR-100 but also exhibits superior transferability, with notable improvements of 3.9% and 4.7% in STL-10 and TIN-200, respectively.
@@
Table 3: Test top-1 accuracy (%) of WRN-16-2 (student) distilled from WRN-40-2 (teacher). In this setup, the representations learned from the CIFAR-100 dataset are transferred to the STL-10 and TIN-200 datasets. The network is frozen, and a linear classifier is trained on the last feature layer to perform classification with 10 classes (STL-10) or 200 classes (TIN-200). Results adapted from [29]. The values in bold indicate the maximum of each row.
@@ 5 Conclusions
We have developed a new technique for neural network distillation that leverages invariant consistency to improve the traditional KD process. Our method, demonstrated for model compression, can also be applied to cross-modal knowledge transfer and ensemble distillation from a group of teachers to a single student network. It combines contrastive learning with an invariance penalty to better align teacher and student models. Unlike CRD [29], our approach requires no memory bank. Instead, it uses a trainable temperature parameter and a bias to scale and shift the logits. By leveraging the contrastive nature, ICD not only aligns the student with the teacher’s output but also improves the quality of the learned embeddings, leading to better generalization and performance. Our experiments demonstrate that this approach consistently outperforms standard KD methods, enhancing both accuracy and robustness. Notably, our method has shown excellent performance on CIFAR-100 and has been successfully applied to other datasets such as TIN-200 and STL-10. In some cases, student models using our KD technique even exceeded the performance of their teacher models, suggesting significant potential for this approach in optimizing neural network training and deployment in various settings.

### v2

@@ 4 Experiments
We evaluate our ICD framework in the KD task for model compression of a large network to a smaller one, similar to [22, 41, 54, 48, 39, 1, 36, 37, 21, 25, 53, 24, 47]. This method aligns with common practices in the field, ensuring a fair comparison between different techniques.
Datasets. (1) CIFAR-100 [26] contains 50,000 training images with 500 images per class and 10,000 test images. (2) ImageNet ILSVRC-2012 [12] includes 1.2 million images from 1,000 classes for training and 50,000 for validation. (3) STL-10 [11] consists of a training set of 5,000 labeled images from 10 classes and 100,000 unlabeled images, and a test set of 8,000 images. (4) Tiny ImageNet (TIN-200) [12] comprises 200 classes, each with 500 training images and 50 validaton images.
Setup. We experiment with student-teacher combinations of different capacities, such as ResNet [19] or Wide ResNet (WRN) [55], VGG [45], MobileNet [42], and ShuffleNet [56, 30] (more details are described in the supplementary material). The only hyperparameters in our method are: the loss coefficient $\beta$, which balances the ICD loss with other loss terms and the internal loss coefficient $\alpha$, which balances the contrastive and invariance losses within the ICD loss. We set $\alpha=0.5$ and $\beta=1$ and ablate both $\alpha$ and $\beta$ in Section 4.5. The hyperparameter $\lambda$ is set to $1.0$ for the KL divergence loss to maintain consistency with [47, 5, 4]. Both the student and teacher outputs are projected to a 128-dimensional space using a projection head consisting of a single linear layer, followed by $\ell_{2}$ normalization. We train for a total of 240 epochs. More details on the training procedures can be found in the supplementary material.
Comparison. We compare our approach to the following state-of-the-art methods: (1) KD [22]; (2) FitNets [41]; (3) AT [54]; (4) SP [48]; (5) CC [39]; (6) VID [1]; (7) RKD [36]; (8) PKT [37]; (9) AB [21]; (10) FT [25]; (11) FSP [53]; (12) NST [24]; (13) CRD [47]; (14) OFD [20]; (15) WSLD [57]; (16) IPWD [35]. In the supplementary material, we include additional methods: (17) SRRL [52]; (18) SemCKD [5]; (19) ReviewKD [8]; (20) SimKD [4]; (21) DistPro [13]; (22) NORM [28]; (23) ITRD [32]; (24) FKD [17]; (25) CRCD [58]; (26) EGA [31]; (27) WCoRD [7].
@@ 4.1 Results on CIFAR-100
Table 1 and Table 2 provide a comprehensive comparison of top-1 accuracies across various KD methods for both identical and differing architectures between student and teacher models on the CIFAR-100 dataset. Specifically, Table 1 focuses on scenarios where the student and teacher share the same architecture, while Table 2 explores settings with differing architectures. Our proposed method, ICD, and its combination with KD, consistently achieve superior performance compared to other distillation objectives, including the original KD. Our method surpasses the teacher network’s performance in both same-architecture (WRN-40-2 to WRN-16-2) and cross-architecture (WRN-40-2 to ShuffleNet-v1) scenarios, achieving accuracy gains of 0.45% and 0.90%, respectively. While CRD and CRD combined with KD also show competitive results, especially in some architecture configurations, ICD and ICD combined with KD frequently show higher accuracies. Both CRD and ICD leverage contrastive learning objectives, which promote better alignment between student and teacher models.
@@
OFD [20] 75.24 ($\uparrow$) 74.33 ($\uparrow$) 70.38 ($\uparrow$) n/a 73.23 ($\uparrow$) 74.95 ($\uparrow$) 73.95 ($\uparrow$)
WSLD [57] n/a 73.74 ($\uparrow$) 71.53 ($\uparrow$) n/a 73.36 ($\uparrow$) 74.79 ($\uparrow$) n/a
IPWD [35] n/a 74.64 ($\uparrow$) 71.32 ($\uparrow$) n/a 73.91 ($\uparrow$) 76.03 ($\uparrow$) n/a
ICD (ours) 74.99 ($\uparrow$) 73.69 ($\uparrow$) 71.18 ($\uparrow$) 71.00 ($\uparrow$) 73.12 ($\uparrow$) 74.23 ($\uparrow$) 73.22 ($\uparrow$)
ICD+KD (ours) 76.06 ($\uparrow$) 74.76 ($\uparrow$) 71.81 ($\uparrow$) 72.03 ($\uparrow$) 73.62 ($\uparrow$) 75.09 ($\uparrow$) 73.95 ($\uparrow$)
Table 1: Test top-1 accuracy (%) of student networks on CIFAR-100, comparing students and teachers of the same architecture using various distillation methods. $\uparrow$ denotes outperformance over KD and $\downarrow$ denotes underperformance. The values in bold indicate the maximum of each column while underlined values mark the second best.
@@
OFD [20] 69.48 ($\uparrow$) 69.04 ($\uparrow$) n/a 75.98 ($\uparrow$) 76.82 ($\uparrow$) 75.85 ($\uparrow$)
WSLD [57] n/a 68.79 ($\uparrow$) 73.80 ($\uparrow$) 75.09 ($\uparrow$) n/a 75.23 ($\uparrow$)
IPWD [35] n/a 70.25 ($\uparrow$) 74.97 ($\uparrow$) 76.03 ($\uparrow$) n/a 76.44 ($\uparrow$)
ICD (ours) 68.35 ($\uparrow$) 67.39 ($\uparrow$) 73.85 ($\uparrow$) 74.26 ($\uparrow$) 75.26 ($\uparrow$) 74.98 ($\uparrow$)
ICD+KD (ours) 69.55 ($\uparrow$) 69.91 ($\uparrow$) 74.08 ($\uparrow$) 75.27 ($\uparrow$) 76.58 ($\uparrow$) 76.51 ($\uparrow$)
Table 2: Test top-1 accuracy (%) of student networks on CIFAR-100 involving students and teachers from different architectures, using various distillation methods. $\uparrow$ denotes outperformance over KD and $\downarrow$ denotes underperformance. The values in bold indicate the maximum of each column while underlined values mark the second best.
4.2 Results on ImageNet
Table 3 showcases the top-1 accuracy of student networks that were trained using various distillation methods on ImageNet. The results demonstrate the effectiveness of our method on large-scale datasets, highlighting its ability to distill knowledge from complex models and enhance the performance of student networks. Our approach achieves competitive results, surpassing KD and attaining state-of-the-art performance in ResNet-50 to ResNet-18 distillation. Our approach also shows improvement across different architectures, demonstrating its effectiveness in various distillation scenarios.
Teacher ResNet-34 ResNet-50 ResNet-50
Student ResNet-18 ResNet-18 MobileNet
Teacher 73.31 76.16 76.16
Student 69.75 69.75 69.63
KD [22] 70.67 71.29 70.49
AT [54] 71.03 ($\uparrow$) 71.18 ($\downarrow$) 70.18 ($\downarrow$)
SP [48] 70.62 ($\uparrow$) 71.08 ($\downarrow$) n/a
CC [39] 69.96 ($\downarrow$) n/a n/a
VID [1] n/a 71.11 ($\downarrow$) n/a
RKD [36] 70.40 ($\downarrow$) n/a 68.50 ($\downarrow$)
AB [21] n/a n/a 68.89 ($\downarrow$)
FT [25] n/a n/a 69.88 ($\downarrow$)
FSP [53] 70.58 ($\downarrow$) n/a n/a
NST [24] 70.29 ($\downarrow$) n/a n/a
CRD [47] 71.17 ($\uparrow$) 71.25 ($\downarrow$) 69.07 ($\downarrow$)
OFD [20] 71.03 ($\uparrow$) n/a 71.33 ($\uparrow$)
WSLD [57] 72.04 ($\uparrow$) n/a 71.52 ($\uparrow$)
IPWD [35] 71.88 ($\uparrow$) n/a 72.65 ($\uparrow$)
ICD (ours) 71.10 ($\uparrow$) 71.38 ($\uparrow$) 70.51 ($\uparrow$)
ICD+KD (ours) 71.71 ($\uparrow$) 71.65 ($\uparrow$) 71.55 ($\uparrow$)
Table 3: Test top-1 (%) on ImageNet validation test using various distillation methods. The table compares students and teachers of the same and different architecture. $\uparrow$ denotes outperformance over KD and $\downarrow$ denotes underperformance. The values in bold indicate the maximum of each column while underlined values mark the second best.
@@ 4.4 Transferability of Representations
Our research focuses on transferring knowledge from a teacher network to a student network while learning representations that embody general knowledge applicable across various tasks and datasets. To investigate this, we employ a distillation process where a WRN-40-2 teacher network transfers its learned representations to a WRN-16-2 student network, which can either be trained directly from the CIFAR-100 dataset or through distillation. In our experiments, the student network functions as a fixed feature extractor, processing images from STL-10 and TIN-200, both resized to $32\times 32$. To assess the generalizability of these representations, we train a linear classifier on top of the last feature layer to perform 10-way classification for STL-10 and 200-way classification for TIN-200. The effectiveness of different distillation methods in enhancing the transferability of these representations is detailed in Table 4. Our results indicate that the ICD method, both standalone and in combination with KD, significantly outperforms all other distillation methods in enhancing the transferability of learned representations across different datasets. This superior transferability suggests that ICD encourages the student to learn more general and robust features, which are less overfitted to the specific training dataset and more applicable to diverse visual recognition tasks.
@@
Table 4: Test top-1 accuracy (%) of WRN-16-2 (student) distilled from WRN-40-2 (teacher). In this setup, the representations learned from the CIFAR-100 dataset are transferred to the STL-10 and TIN-200 datasets. The network is frozen, and a linear classifier is trained on the last feature layer to perform classification with 10 classes (STL-10) or 200 classes (TIN-200). The values in bold indicate the maximum of each row.
4.5 Ablation Study
Figure 2: Ablation study results on CIFAR-100 using WRN-40-2 as the teacher and WRN-16-2 as the student. (a) Effect of the internal ICD coefficient $\alpha$ on performance. (b) Effect of ICD loss coefficient $\beta$ on performance (logarithmic scale). (c) Effect of loss coefficient $\lambda$ on performance.
There are three main hyperparameters in our objective: the internal ICD coefficient $\alpha$, which balances the contrastive and invariance losses within the ICD loss; the ICD loss coefficient $\beta$, which balances the ICD loss with other loss terms; and the loss coefficient $\lambda$, which is typically set to 1.0 but can be adjusted to affect the weighting of certain components. We conduct an ablation study to analyze the impact of these hyperparameters. For this study, we adopt WRN-40-2 as the teacher and WRN-16-2 as the student. Experiments are conducted on CIFAR-100, and the results are shown in Figure 2.
Internal loss coefficient $\bm{\alpha}$. We tested different values for $\alpha$: 0.01, 0.1, 0.3, 0.5, 0.7, 1, 2, and 5. As shown in Figure 2(a), very low $\alpha$ values degrade performance, while increasing $\alpha$ beyond 1 slightly reduces performance. This suggests that a balanced contribution from both the contrastive and invariance components of the ICD loss is crucial for optimal KD.
Loss coefficient $\bm{\beta}$. We varied $\beta$ from 0.1 to 100, considering values of 0.1, 0.5, 1, 2, 5, 10, 50, and 100. As illustrated in Figure 2(b), extremely high $\beta$ values cause significant degradation in performance due to the overwhelming contribution of the ICD loss compared to other loss terms. Very low values of $\beta$ also lead to decreased performance. The optimal range for $\beta$ is between 0.5 and 10, suggesting that the ICD loss should be weighted similarly to other loss terms for the best results.
Loss coefficient $\bm{\lambda}$. While $\lambda$ is typically set to 1.0 [47, 5, 4], we tested values from 0.1 to 100. As shown in Figure 2(c), high values (i.e., $\lambda=50$ and $100$) lead to collapsing training. Lower values, such as 0.5 to 1.0, have similar performance.
@@ 5 Conclusions
We have developed a new technique for neural network distillation that leverages invariant consistency to improve the traditional KD process. Our method, demonstrated for model compression, can also be applied to cross-modal knowledge transfer and ensemble distillation from a group of teachers to a single student network. It combines contrastive learning with an invariance penalty to better align teacher and student models. Unlike CRD [47], our approach requires no memory bank. Instead, it uses a trainable temperature parameter and a bias to scale and shift the logits. By leveraging the contrastive nature, ICD not only aligns the student with the teacher’s output but also improves the quality of the learned embeddings, leading to better generalization and performance. Our experiments demonstrate that this approach consistently outperforms standard KD methods, enhancing both accuracy and robustness. Notably, our method has shown excellent performance on CIFAR-100 and ImageNet and has been successfully applied to other datasets such as TIN-200 and STL-10. In some cases, student models using our KD technique even exceeded the performance of their teacher models, suggesting significant potential for this approach in optimizing neural network training and deployment in various settings.